OpenStudy (anonymous):

USE THE DISTRIBUTIVE PROPERTY TO REWRITE EACH EXPRESSION. 6x(y-z)

OpenStudy (anonymous):

6xy-6xz

OpenStudy (anonymous):

thanks wb -3b(2b-2a)

OpenStudy (anonymous):

-6b^2 + 6 ab

OpenStudy (anonymous):

DOES THAT ^2 THAT LIL AROW THIN IS THAT TO THE POWER?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

OKAA...:)

OpenStudy (anonymous):

THNKS :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

WB M(2M+1)

OpenStudy (anonymous):

2M^2 + M

OpenStudy (anonymous):

R U A 100% SURE:)?

OpenStudy (anonymous):

yeah :) 10000% sure

OpenStudy (anonymous):

lol <3 :) wb -7(5-n)

OpenStudy (anonymous):

-35+7n

OpenStudy (anonymous):

i thouight u subtract not add

OpenStudy (anonymous):

It's because there are two negatives -7 and - n so it becomes a positive

OpenStudy (anonymous):

okaa.. thank u soooooooo muchh

OpenStudy (anonymous):

:) ur welcome

OpenStudy (anonymous):

:)<3<3

OpenStudy (anonymous):

now can u evaluate 5 7/9+ 2 1/4 = idk help lol

OpenStudy (anonymous):

okay wait :)

OpenStudy (anonymous):

okayy...:)

OpenStudy (anonymous):

It would be 8 1/36

OpenStudy (anonymous):

can u show me how u got it?

OpenStudy (anonymous):

can u show ur work plzz?

OpenStudy (anonymous):

okay 5 7/9+ 2 1/4 You have to find the LCM of the denominators 9 and 4 which is 36 so then you make complete set of fractions 5 7/9 = 5 28/36 2 1/4 = 2 9/36 when you add you get 7 37/36 But then you simplify 37/36 which is 1 1/36 so that plus 7 equals 8 1/36

OpenStudy (anonymous):

okaa.... thnks im srry that im askin all these questions i have more but i dont know if u want me too keep askin u allthese questions

OpenStudy (anonymous):

:) iT'S OKAY cmon what u have?

OpenStudy (anonymous):

okaaa r u sure:)?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

oka boo boo:)

OpenStudy (anonymous):

7 1/4-4 5/6=

OpenStudy (anonymous):

Okay so first you have to map the fractions into improper fractions, so 7 1/4 = 29/4 4 5/6 = 29/6 Then you Find the LCM which is 12 and make them new fractions 29/4 = 87/12 29/6 = 58/12 When you subtract you get 29/12 when you make that into a complex fraction which is 2 5/12